# CTP-SCU/2025004

### v2

@@ Abstract
We present an operator-algebraic definition for timelike entanglement entropy in QFT. This rigorously defined timelike entanglement entropy is real-valued due to the timelike tube theorem.
@@ Introduction
$S(\mathcal{O}):=\lim_{\widetilde{\mathcal{O}}\rightarrow\mathcal{\mathcal{O}}}S_{\mathcal{A}}(\mathcal{O},\widetilde{\mathcal{O}}).$ (5)
@@
where $\beta$ is the inverse temperature. For a finite-size CFT2 on ×$\mathbb{R}\times S^{1}$, the timelike entanglement entropy becomes
@@
One important question is how to understand within the Euclidean path integral approach, that the timelike entanglement entropy defined here is real-valued, given that the pure state timelike entanglement entropy introduced in [3] and subsequent works is complex-valued. To this end, let us consider the density matrix shown in Figure 2, on which the entanglement entropy between disjoint symmetric spacelike intervals $A$ and $B$ is computed [17, 18],
@@
The usual pure state is a limit of this disjoint system. It is evident that the path integrals on this single annulus $Z_{1}$ and replica $n-$folded annulus $Z_{n}$ are both independent of specific intervals $A$ and $B$. Consequently, the entanglement entropy between any two disjoint segments on this annulus is always (15). The asymmetric configurations are obtained by conformal transformations which do not give rise to complex numbers. Thus, for any disjoint pairwise combination of timelike, spacelike, and lightlike regions, the entanglement entropy $S_{\rm vN}(A:B)$ is always real-valued. The pure state half-space entanglement entropy is achieved by setting $R_{1}\to\epsilon$ and taking $R_{2}$ as the IR cutoff in equation (15). Note the pure state entanglement entropy $S_{\rm vN}=c/3\log\ell/\epsilon$ for a finite length $\ell$ is obtained from $S_{\rm vN}(C:D)$ in the proper limits [17, 18].
Figure 2: The density matrix for disjoint symmetric segments $A$ and $B$.
@@ Acknowledgments.
We deeply appreciate Edward Witten for pointing out that the strong subadditivity fails by the non-commutativity of the algebras in the timelike direction. This work is supported in part by NSFC (Grant No. 12105191 and 12275184).

### v3

@@ Abstract
We present an operator-algebraic definition for timelike entanglement entropy in CFT. This rigorously defined timelike entanglement entropy is real-valued due to the timelike tube theorem. We further demonstrate why the timelike entanglement entropy should be real-valued from both path integral argument and holography perspective.
@@ Introduction
[…]
@@
where $\beta$ is the inverse temperature. For a finite-size CFT2 on $\mathbb{R}\times S^{1}$, the timelike entanglement entropy becomes
@@
Why should timelike entanglement entropy be real-valued?
Recent literature has predominantly considered a complex-valued timelike entanglement entropy, while our proposal yields a real-valued result. It is therefore important to resolve this contradiction. We work this out from both path integral argument and holography perspective.
.1 Path integral argument
Let us consider the density matrix shown in Figure 2, on which the entanglement entropy between disjoint symmetric spacelike intervals $A$ and $B$ is computed [17, 18],
@@
The usual divergent entanglement entropy between adjacent regions is a simple adjacent limit of this disjoint system. It is evident that the path integrals on this single annulus $Z_{1}$ and replica $n$-folded annulus $Z_{n}$ are both independent of specific intervals $A$ and $B$. Consequently, the entanglement entropy between any two disjoint segments on this annulus is always equation (15). The asymmetric configurations are obtained by conformal transformations which do not give rise to complex numbers. Thus, for any disjoint pairwise combination of timelike, spacelike, and lightlike regions, the entanglement entropy $S_{\rm vN}(A:B)$ is always real-valued. The half-space entanglement entropy is achieved by setting $R_{1}\to\epsilon$ and taking $R_{2}\to\xi$ as the IR cutoff in equation (15). Note the usual divergent adjacent entanglement entropy $S_{\rm vN}=c/3\log\ell/\epsilon$ for a finite length $\ell$ is obtained from $S_{\rm vN}(C:D)$ in the proper limits [17, 18]. From this picture, it is clear that if one only Wick rotates $R_{2}$ (i.e. $\ell$) but not simultaneously rotates $R_{1}$ (i.e. $\epsilon$), a spurious complex number appears.
Figure 2: The density matrix for disjoint symmetric segments AA and $B$. The usual adjacent configurations with divergent entanglement entropies $c/3\log\ell/\epsilon$ or $c/6\log\xi/\epsilon$ are obtained by taking simple limits.
.2 Holography perspective
Without loss of generality, consider the entanglement entropy of a single timelike interval $\mathcal{T}_{1}=\left\{0<t<1,x=0\right\}$ in zero-temperature CFT2. This interval can be mapped onto the temporal half-axis $\mathcal{T}_{\text{h}}=\left\{t>0,x=0\right\}$ via a special conformal transformation,
$x^{\mu}\rightarrow\frac{x^{\mu}-b^{\mu}x^{2}}{1-2(b\cdot x)+b^{2}x^{2}},$ (16)
with a timelike vector $b^{\mu}=(1,0)$. Therefore, the entanglement entropy of any finite timelike interval $\mathcal{T}$ equals that of the temporal half-axis $\mathcal{T}_{\text{h}}$ due to conformal symmetry:
$S(\mathcal{T})=S(\mathcal{T}_{\text{h}}).$ (17)
Figure 3: Penrose diagrams of Poincaré patch of AdS3 spacetime and holographic duals of $S(\mathcal{T})$, $S(\mathcal{T}_{\text{h}})$. In the left panel, the blue line represents a finite timelike interval $\mathcal{T}$ on the conformal boundary of AdS3, the red line denotes a timelike geodesic, and the two green lines denote spacelike geodesics. In the right panel, the blue line represents the temporal half-axis $\mathcal{T}_{\text{h}}$, and the green line denotes a spacelike geodesic.
It has been conjectured that the holographic dual of $S(\mathcal{T})$ consists of two spacelike geodesics and one timelike geodesic connecting them in AdS3 spacetime, with the metric
$\mathrm{d}s^{2}=\frac{\mathrm{d}z^{2}+\mathrm{d}x^{\mu}\mathrm{d}x_{\mu}}{z^{2}},$ (18)
as illustrated in Figure 3. The timelike geodesic is argued to contribute an imaginary part to $S(\mathcal{T})$.
However, this conjecture fails for $S(\mathcal{T}_{\text{h}})$, whose holographic dual is a single spacelike geodesic from $(z,x^{\mu})=(0,0)$ to $(z,x^{\mu})=(+\infty,0)$, containing no timelike geodesics. Thus, from the holographic perspective, $S(\mathcal{T}_{\text{h}})$ is purely real-valued, and by the equivalence in (17), $S(\mathcal{T})$ must also be real-valued.
@@ Acknowledgments.
We deeply appreciate Edward Witten for pointing out that the strong subadditivity does not hold on the timelike direction by the non-commutativity of the algebras. This fails the simple proof of $F$-theorem included in the first version of this paper. This work is supported in part by NSFC (Grant No. 12105191 and 12275184).